### v1

# Oblique boundary value problems for degenerate Hessian quotient type equations

Abstract.
In this article, we study oblique boundary value problems for degenerate Hessian quotient type equations in bounded domains without any geometric conditions. By establishing the a priori estimates, we show the existence and uniqueness of $C^{1,1}$ admissible solutions.
Key words and phrases:
Degenerate Hessian quotient type equations, Oblique boundary value problems.
2020 Mathematics Subject Classification:
35J70, 35J66.
This research was supported by funds from the National Natural Science Foundation of China No. 11971157.
$\ast$ Corresponding author
1. Introduction
In this article, we consider oblique boundary value problems for degenerate Hessian quotient type equations
(1.1) $&\frac{\sigma_{k}(\gamma\triangle uI-D^{2}u)}{\sigma_{l}(\gamma\triangle uI-D^{2}u)}=f(x,u),\,\,\,\text{in}\,\;\Omega,\\ &G(x,u,Du)=0,\,\,\,\,\quad\quad\quad\quad\quad\text{on}\,\;\partial\Omega,\end{aligned}\right.$
where $\Omega$ is a bounded domain in $\mathbb{R}^{n}$, $\gamma\geq 1$, $0\leq l<k\leq n$, $Du$ and $D^{2}u$ denote the gradient vector and the Hessian matrix of the solution $u$. For convenience, let $x,z,p$ be points in $\Omega,\mathbb{R},\mathbb{R}^{n}$, respectively. Set $f(x,z)$ is a non-negative smooth function defined on $\Omega\times\mathbb{R}$. The boundary value problem $G(x,z,p)\in C^{1}(\partial\Omega\times\mathbb{R}\times\mathbb{R}^{n})$ is said to be oblique with respect to $u$, if
(1.2) $G_{p}\cdot\nu\geq\beta_{0},\,\,\text{for all}\,\;(x,z,p)\in\partial\Omega\times\mathbb{R}\times\mathbb{R}^{n},$
where $\nu$ is the unit inner normal vector field on $\partial\Omega$ and $\beta_{0}$ is a positive constant. If $G_{p}\cdot\nu>0$ on all of $\partial\Omega\times\mathbb{R}\times\mathbb{R}^{n}$, we will simply refer to $G$ as oblique. As in Jiang-Trudinger-Xiang [24], we will assume that the function $G\in C^{2}(\partial\Omega\times\mathbb{R}\times\mathbb{R}^{n})$ is concave in $p$, that is $G_{pp}\leq 0$ on $\partial\Omega\times\mathbb{R}\times\mathbb{R}^{n}$. This includes the quasilinear case, when $G_{pp}=0,$
(1.3) $G(x,z,p)=\beta(x,z)\cdot p-\varphi(x,z)=0,\quad\text{on}\;\partial\Omega\times\mathbb{R}\times\mathbb{R}^{n},$
where $\beta=G_{p}$ and $\varphi$ are defined on $\partial\Omega\times\mathbb{R}$. In this article, we focus on the semilinear case,
(1.4) $u_{\beta}=\varphi(x,u),\quad\text{on}\ \partial\Omega\times\mathbb{R},$
where $\beta\cdot\nu>0$. Especially, we need $\varphi_{u}\geq\gamma_{0}>0$ which is only used to prove upper bounds of the admissible solution.
Next, we recall some definitions about Hessian equations.
Definition 1.1.
Let $\lambda=(\lambda_{1},\dots,\lambda_{n})\in\mathbb{R}^{n}$, we set the elementary symmetric function for $k=1,2,\cdots,n$
$\sigma_{k}(\lambda)=\sum_{1\leq i_{1}<i_{2}<\cdots<i_{k}\leq n}\lambda_{i_{1}}\lambda_{i_{2}}\cdots\lambda_{i_{k}},$
specially, denote $\sigma_{0}=1$ and $\sigma_{k}=0$ for $k>n$ or $k<0$.
Definition 1.2.
[…]
Hessian quotient type equations
(1.5) $\frac{\sigma_{k}(\gamma\triangle uI-D^{2}u)}{\sigma_{l}(\gamma\triangle uI-D^{2}u)}=f$
originate from many geometric problems and play an important role in conformal geometry, complex geometry. Equation (1.5) is called the $(n-1)$ Monge-Ampère equation which was introduced by Harvey-Lawson in a series of papers [16, 17, 18] when $k=n$, $l=0$ and $\gamma=1$. And more information can be consulted in [10, 8, 9, 14, 35, 40, 41] and references therein.
For non-degenerate cases, Dirichlet problems for basic Hessian equations have been well studied by Caffarelli-Nirenberg-Spruck [1], Ivochkina [19] and Trudinger [39]. Some works about Neumann problems for basic Hessian equations can be found in [4, 33, 36, 38]. Dong-Wei [7] have established the a priori estimates to equation (1.5) when $\lambda(\triangle uI-D^{2}u)(x)\in\Gamma_{k+1}$. Then Chen-Dong-Han [5] showed that the existence and uniqueness of the admissible solution to equation (1.5) with $0\leq l<k<n$ when $\lambda(\triangle uI-D^{2}u)(x)\in\Gamma_{k}$.
Let’s review some research about degenerate fully nonlinear equations. The a priori $C^{1,1}$ estimates play an important role in the proof of $C^{1,1}$ solutions in degenerate problems. Guan [12] established $C^{1,1}$ estimates to Dirichlet problems of degenerate Monge-Ampère equations for homogenous boundary condition with $f^{\frac{1}{n}}\in C^{1,1}(\overline{\Omega})$, and Guan-Trudinger-Wang [13] for non-homogenous boundary condition with $f^{\frac{1}{n-1}}\in C^{1,1}(\overline{\Omega})$. Dong [6] obtained $C^{1,1}$ estimates to Dirichlet problems of degenerate Hessian equations for homogenous boundary condition with $f^{\frac{1}{k-1}}\in C^{1,1}(\overline{\Omega})$ and $Df^{\frac{1}{k-1}}\leq Cf^{\frac{1}{2(k-1)}}$. Recently, Jiao-Wang [20] solved $C^{1,1}$ estimates to convex solutions of Dirichlet problems for degenerate Hessian equations with non-homogenous boundary conditions defined in the strictly convex bounded domain with $f^{\frac{1}{k-1}}\in C^{1,1}(\overline{\Omega})$. If $f^{\frac{1}{k}}\in C^{1,1}(\overline{\Omega})$ for non-homogenous boundary problems, $C^{1,1}$ estimates have been obtained by Krylov [26, 27, 28, 29]. Wang [47] gave an example to show that the condition $f^{\frac{1}{k-1}}\in C^{1,1}(\overline{\Omega})$ is sharp for degenerate Hessian equations. The existence of $C^{1,1}$ solutions to Neumann problems for degenerate basic Hessian quotient equations with $f^{\frac{1}{k-l}}\in C^{1,1}(\overline{\Omega})$ was proved by Mei [34]. For more reference, the reader may see [2, 11, 48] and the reference therein.
Therefore, it is of natural interest to consider existence results for the oblique boundary value problem. The oblique boundary value problem for Hessian type equations arises naturally in the theory of fully nonlinear elliptic equations, through its applications in conformal geometry, optimal transportation, and geometric optics. Wang [45] derived the existence of solutions for two dimensional Monge-Ampère equations of the oblique boundary value problem and considered the generalized solutions for general dimension. The oblique boundary value problem of linear and quasilinear elliptic equations can be seen in the book [30]. Lieberman-Trudinger [31] considered the nonlinear oblique boundary value problem for quasilinear and fully nonlinear uniformly elliptic partial differential equations under some structure conditions. Urbas [42, 43, 44] studied Monge-Ampère equations, two dimensional Hessian equations and curvature equations of the oblique boundary value problem. For augmented Hessian equations with the oblique boundary value problem
$\begin{cases}F(D^{2}u-A(x,z,p))=B(x,z,p),\,\,\,\text{in}\,\;\Omega,\\ G(x,z,p)=0,\,\,\,\,\quad\quad\quad\quad\quad\quad\quad\quad\text{on}\,\;\partial\Omega,\end{cases}$
a series works of Jiang-Trudinger [21, 22, 23] considered existence results by formulating some conditions on $A$ and $F$, which provided a comprehensive framework for studying the oblique boundary value problem for a class of fully nonlinear equations. Then Wang [46] derived global gradient estimates for admissible solutions to Hessian equations in $\mathbb{R}^{n}$ with the oblique boundary value problem.
Inspired by the above works, we naturally consider the oblique boundary value problem (1.1) for degenerate Hessian quotient type equations and get the following main theorem. In order to obtain the existence, we need to assume the existence of an admissible subsolution $\underline{u}$ to problem (1.1) satisfying
(1.6) $&\frac{\sigma_{k}(\gamma\triangle\underline{u}I-D^{2}\underline{u})}{\sigma_{l}(\gamma\triangle\underline{u}I-D^{2}\underline{u})}\geq f(x,\underline{u}),\,\,\,\text{in}\,\;\Omega,\\ &G(x,\underline{u},D\underline{u})=0,\,\,\,\,\quad\quad\quad\quad\quad\text{on}\,\;\partial\Omega.\end{aligned}\right.$
Then we have the following existence result for $\gamma\geq 1$ and $0\leq l<k<n$.
Theorem 1.3.
Let $u\in C^{4}(\Omega)\cap C^{3}(\overline{\Omega})$ be an admissible solution of problem (1.1) in a bounded $C^{4}$ domain $\Omega\subset\mathbb{R}^{n}$, $f^{\frac{1}{k-l}}(x,z)\in C^{1,1}(\overline{\Omega}\times\mathbb{R})$ and $f_{z}\geq 0$. Suppose that $G\in C^{2}(\partial\Omega\times\mathbb{R}\times\mathbb{R}^{n})$ satisfy the oblique boundary condition (1.4). Assume that $\underline{u}$ is an admissible subsolution of problem (1.1). Then there exists a unique solution $u\in C^{1,1}(\overline{\Omega})$ of problem (1.1) satisfying
(1.7) $|u|_{C^{1,1}(\overline{\Omega})}\leq C,$
where the positive constant $C$ depends on $n,\ k,\ l,\ \gamma_{0},\ \gamma,\ \overline{\Omega},$ $\underline{u}$, $G$, $|f^{\frac{1}{k-l}}|_{C^{1,1}}$.
Remark 1.4.
The assumption of the admissible subsolution and $f_{z}\geq 0$ are only used to prove maximum modulus estimates. Furtherly, When $\gamma>1$ in problem (1.1), the above existence result still holds for $k=n$.
Remark 1.5.
When the matrix $\left\{\gamma\triangle uI-D^{2}u\right\}$ ($\gamma\geq 1$) is replaced by $\left\{\gamma\triangle uI+D^{2}u\right\}$ ($\gamma>0$) in problem (1.1), the existence theorem still holds for the problem
(1.8) $&\frac{\sigma_{k}(\gamma\triangle uI+D^{2}u)}{\sigma_{l}(\gamma\triangle uI+D^{2}u)}=f(x,u),\,\,\,\text{in}\,\;\Omega,\\ &G(x,u,Du)=0,\quad\quad\quad\quad\quad\ \ \text{on}\,\;\partial\Omega,\end{aligned}\right.$
where $0\leq l<k\leq n$.
The rest of the paper is organized as follows. In Section 2, some preliminaries are provided. Gradient estimates are established in Section 3. Section 4 are devoted to second order estimates and the proof of Theorem 1.3.
2. Preliminaries
In this section, we recall some basic properties of Hessian quotient type equations, which are well-known and can be found in [3, 5, 7].
Firstly, we define some notations for convenience. Let $U=\gamma\triangle uI-D^{2}u,U_{ij}=\gamma\triangle u\delta_{ij}-u_{ij}$, $\lambda=(\lambda_{1},\lambda_{2},\cdots,\lambda_{n})$ and $\eta=(\eta_{1},\eta_{2},\cdots,\eta_{n})$ are eigenvalue vectors of matrix $D^{2}u$ and $U$, respectively. Also, we set
(2.1) $F(u_{ij})=\frac{\sigma_{k}(U_{ij})}{\sigma_{l}(U_{ij})},\quad F^{ij}=\frac{\partial F}{\partial u_{ij}},\quad F^{ij,rs}=\frac{\partial^{2}F}{\partial u_{ij}\partial u_{rs}},$
and
(2.2) $\widetilde{F}(u_{ij})=\left[\frac{\sigma_{k}(U_{ij})}{\sigma_{l}(U_{ij})}\right]^{\frac{1}{k-l}},\quad\widetilde{F}^{ij}=\frac{\partial\widetilde{F}}{\partial u_{ij}},\quad\widetilde{F}^{ij,rs}=\frac{\partial^{2}\widetilde{F}}{\partial u_{ij}\partial u_{rs}},$
thus Hessian quotient type equations in problem (1.1) can be written as
(2.3) $\widetilde{F}(u_{ij})=F^{\frac{1}{k-l}}(u_{ij})=\left[\frac{\sigma_{k}(U_{ij})}{\sigma_{l}(U_{ij})}\right]^{\frac{1}{k-l}}=f^{\frac{1}{k-l}}(x,u)=\widetilde{f}(x,u).$
Differentiating equation (2.3) with respect to vector fields $\xi=(\xi^{1},\xi^{2},\cdots,\xi^{n})$ twice with $\xi^{i}\in C^{2}(\overline{\Omega})$, $i=1,2,\cdots,n.$, we have
(2.4) $F^{ij}u_{ij\xi}=\partial_{\xi}f:=f_{x_{\xi}}+f_{z}u_{\xi},$
(2.5) $\widetilde{F}^{ij}u_{ij\xi}=\partial_{\xi}\widetilde{f}:=\widetilde{f}_{x_{\xi}}+\widetilde{f}_{z}u_{\xi},$
and
(2.6) $F^{ij,rs}u_{ij\xi}u_{rs\xi}+F^{ij}u_{ij\xi\xi}=\partial_{\xi\xi}f:=f_{x_{\xi}x_{\xi}}+2f_{x_{\xi}z}u_{\xi}+f_{zz}u_{\xi}^{2}+f_{z}u_{\xi\xi},$
(2.7) $\widetilde{F}^{ij,rs}u_{ij\xi}u_{rs\xi}+\widetilde{F}^{ij}u_{ij\xi\xi}=\partial_{\xi\xi}\widetilde{f}:=\widetilde{f}_{x_{\xi}x_{\xi}}+2\widetilde{f}_{x_{\xi}z}u_{\xi}+\widetilde{f}_{zz}u_{\xi}^{2}+\widetilde{f}_{z}u_{\xi\xi}.$
Next, we will introduce some properties of elementary symmetric functions.
Definition 2.1.
Let $\lambda=(\lambda_{1},\dots,\lambda_{n})\in\mathbb{R}^{n}$.
(1) $\sigma_{l}(\lambda|i)=\sigma_{l}(\lambda)|_{\lambda_{i}=0},$ $\sigma_{k-1}(\lambda|i)$ is also denoted by $\sigma_{k}^{ii}=\frac{\partial\sigma_{k}}{\partial\lambda_{i}},$
(2) $\sigma_{l}(\lambda|ij)=\sigma_{l}(\lambda)|_{\lambda_{i}=\lambda_{j}=0},$ $\sigma_{k-2}(\lambda|ij)$ is also denoted by $\sigma_{k}^{ii,jj}=\frac{\partial^{2}\sigma_{k}}{\partial\lambda_{i}\partial\lambda_{j}}.$
Lemma 2.2.
Let $\lambda=(\lambda_{1},\dots,\lambda_{n})\in\Gamma_{k}$ and $1\leq k\leq n$, suppose that $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n},$ then
$\sigma_{k-1}(\lambda|k)\geq C(n,k)\sigma_{k-1}(\lambda),\quad\sigma_{k-1}(\lambda|k)\geq C(n,k)\sum_{i=1}^{n}\sigma_{k-1}(\lambda|i),$
where $C(n,k)$ is a positive constant only depending on $n$ and $k$.
The proof of Lemma 2.2 and the following generalized Newton-MacLaurin inequality can be respectively found in [32] and [37].
Lemma 2.3.
For $\lambda\in\Gamma_{k}$, $n\geq k>l\geq 0$, $n\geq r>s\geq 0$, $k\geq r$ and $l\geq s$, we have
$\left[\frac{\sigma_{k}(\lambda)/C_{n}^{k}}{\sigma_{l}(\lambda)/C_{n}^{l}}\right]^{\frac{1}{k-l}}\leq\left[\frac{\sigma_{r}(\lambda)/C_{n}^{r}}{\sigma_{s}(\lambda)/C_{n}^{s}}\right]^{\frac{1}{r-s}},$
and the equality holds if and only if $\lambda_{1}=\cdots=\lambda_{n}>0$.
Lemma 2.4.
If $\lambda=(\lambda_{1},\lambda_{2},\cdots,\lambda_{n})\in\widetilde{\Gamma}_{k}$, then $\left[\frac{\sigma_{k}(\eta)}{\sigma_{l}(\eta)}\right]^{\frac{1}{k-l}}(0\leq l<k\leq n)$ are concave with respect to $\lambda$. Hence, for any $\xi\in\mathbb{R}^{n}$, we have
$\sum_{i,j,r,s}\frac{\partial^{2}\left[\frac{\sigma_{k}(\eta)}{\sigma_{l}(\eta)}\right]}{\partial u_{ij}\partial u_{rs}}\leq\left(1-\frac{1}{k-l}\right)\frac{\left[\sum\limits_{i,j}\frac{\partial\left[\frac{\sigma_{k}(\eta)}{\sigma_{l}(\eta)}\right]}{\partial u_{ij}}\xi_{ij}\right]^{2}}{\frac{\sigma_{k}(\eta)}{\sigma_{l}(\eta)}},$
where $\widetilde{\Gamma}_{k}$ is an open convex symmetric cone in $\mathbb{R}^{n}$, given by
$\widetilde{\Gamma}_{k}=\left\{\lambda\in\mathbb{R}^{n}:\sigma_{i}(\eta)>0,\;\forall\;i=1,\cdots,k\right\}.$
More details of the above lemma can be seen in [5].
Lemma 2.5.
Suppose $\lambda=(\lambda_{1},\lambda_{2},\cdots,\lambda_{n})\in\widetilde{\Gamma}_{k}$ with $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n}$, $\eta_{i}=\gamma\sum_{j=1}^{n}\lambda_{j}-\lambda_{i}$, for $0\leq l<k<n$ we obtain
(1) $\frac{\partial\left[\frac{\sigma_{k}(\eta)}{\sigma_{l}(\eta)}\right]}{\partial\eta_{1}}\geq\frac{\partial\left[\frac{\sigma_{k}(\eta)}{\sigma_{l}(\eta)}\right]}{\partial\eta_{2}}\geq\cdots\geq\frac{\partial\left[\frac{\sigma_{k}(\eta)}{\sigma_{l}(\eta)}\right]}{\partial\eta_{n}},$
(2) $\frac{\partial\left[\frac{\sigma_{k}(\eta)}{\sigma_{l}(\eta)}\right]}{\partial\lambda_{1}}\leq\frac{\partial\left[\frac{\sigma_{k}(\eta)}{\sigma_{l}(\eta)}\right]}{\partial\lambda_{2}}\leq\cdots\leq\frac{\partial\left[\frac{\sigma_{k}(\eta)}{\sigma_{l}(\eta)}\right]}{\partial\lambda_{n}},$
(3) $\frac{\partial\left[\frac{\sigma_{k}(\eta)}{\sigma_{l}(\eta)}\right]}{\partial\lambda_{i}}\geq\frac{2\gamma-1}{n\gamma-1}C(n,k,l)\sum_{j=1}^{n}\frac{\partial\left[\frac{\sigma_{k}(\eta)}{\sigma_{l}(\eta)}\right]}{\partial\lambda_{j}}\geq\frac{2\gamma-1}{n\gamma-1}C(n,k,l)f^{1-\frac{1}{k-l}}\;(1\leq i\leq n).$
Remark 2.6.
When the matrix $\left\{\gamma\triangle uI-D^{2}u\right\}$ ($\gamma\geq 1$) is replaced by $\left\{\gamma\triangle uI+D^{2}u\right\}$ ($\gamma>0$), the similar Lemma 2.4 and Lemma 2.5 still hold for $0\leq l<k\leq n$. Furthermore, gradient estimates and second order estimates can be obtained and we omit details of the proof.
3. Gradient estimates
In this section, we prove gradient estimates for admissible solutions of problem (1.1).
Theorem 3.1.
Let $u\in C^{3}(\Omega)\cap C^{2}(\overline{\Omega})$ be an admissible solution of problem (1.1) in a smooth bounded domain $\Omega\subset\mathbb{R}^{n}$. Assume $f^{\frac{1}{k-l}}(x,z)\in C^{1,1}(\overline{\Omega}\times\mathbb{R})$ and $\sup_{\Omega}|u|\leq M$, where $M$ is a positive constant. If $G(x,z,p)$ is the form of (1.4), suppose that $\beta$ is a smooth unit vector field along $\partial\Omega$ with $\beta,\nu\right>\geq\beta_{0}>0$ for some positive constant $\beta_{0}$. Then, for any $0\leq l<k<n$,
(3.1) $\sup_{\overline{\Omega}}|Du|\leq C,$
where $C>0$ is a constant depending only on $M,n,k,l,\gamma,\overline{\Omega},\beta_{0},|\beta|_{C^{3}},|f^{\frac{1}{k-l}}|_{C^{1}},|\varphi|_{C^{3}}.$
Proof.
Due to Dong-Wei [7], we have already known interior gradient estimates, so we only need to obtain gradient estimates near the boundary. Denote $\Omega_{\mu}=\left\{x\in\Omega\;|\;d(x)\leq\mu<\mu_{0}\leq 1\right\}$, here $\mu,\;\mu_{0}$ are positive constants to be determined later. Then, extend $\beta$ to $\Omega_{\mu}$ smoothly such that $\beta,Dd\right>=\cos\theta\geq\beta_{0}>0$.
Take a unit normal moving frame along $\partial\Omega$, denoted by $\{e_{1},e_{2},\cdots,e_{n-1},\nu\}$, then
(3.2) $\beta=\beta_{\nu}\cdot\nu+\sum_{q=1}^{n-1}\beta_{q}e_{q},$
where $\beta_{\nu}=\left<\beta,\nu\right>=\cos\theta\geq\beta_{0}>0$. Denote by $d(x)=dist(x,\partial\Omega)$ the distance from $x$ to $\partial\Omega$ and let $w(x)=u(x)-\frac{\varphi(x,z)d(x)}{\cos\theta}$. On $\partial\Omega$, since $\varphi(x,z)=u_{\nu}\beta_{\nu}+\sum_{q=1}^{n-1}u_{q}\beta_{q}$, we have
(3.3) $w_{\nu}=-\sum_{q=1}^{n-1}w_{q}\frac{\beta_{q}}{\beta_{\nu}}.$
Applying the fact $|\beta|^{2}=1$ and $\cos\theta>0$,
(3.4) $w_{\nu}^{2}\leq|Dw|^{2}\sin^{2}\theta.$
Consider an auxiliary function
(3.5) $\widetilde{G}(x)=\log\psi(x)+g(u)+\alpha_{1}d(x),$
where
(3.6) $\psi(x)=|Dw|^{2}-\left(\sum_{i=1}^{n}w_{i}d_{i}\right)^{2}=\sum_{i,j=1}^{n}\left(\delta_{ij}-d_{i}d_{j}\right)w_{i}w_{j}=\sum_{i,j=1}^{n}C^{ij}w_{i}w_{j}>0,$
(3.7) $g(u)=\frac{1}{4}\ln\frac{1}{(3M-u)},$
and $\alpha_{1}$ is a large enough positive constant to be determined later. Suppose that $\widetilde{G}(x)$ attains the maximum at $x_{0}\in\overline{\Omega}_{\mu}$, according to interior gradient estimates, we can only consider the following two cases.
$\mathbf{Case\;\mathrm{\mathbf{I}}}$: $x_{0}\in\partial\Omega$. Choosing a coordinate around $x_{0}$ such that $\nu=\frac{\partial}{\partial x_{n}}$ and $\frac{\partial}{\partial x_{s}}\;(s=1,2,\cdots,n-1)$ are tangent to $\partial\Omega$, then
$\frac{\partial d}{\partial x_{n}}=1,\,\frac{\partial d}{\partial x_{s}}=0,\,\frac{\partial^{2}d}{\partial x_{n}\partial x_{i}}=0\,(1\leq i\leq n),\,\frac{\partial^{2}d}{\partial x_{s}\partial x_{q}}=-\kappa_{s}\delta_{sq}\,(1\leq s,q\leq n-1),$
here $\kappa_{s}$ are principal curvatures of $\partial\Omega$ at $x_{0}$. A direct calculation yields
(3.8) $w_{\nu i}=w_{i\nu}+\sum_{j=1}^{n}\mathcal{K}_{ij}w_{j},\quad\text{for}\;i=1,2,\cdots,n,$
where $\mathcal{K}_{ij}$ denoted by the Weingarten matrix of the boundary with respect to $\nu$. In order to derive the later calculations, we deal with $\psi_{\nu}$ in advance
(3.9) $\psi_{\nu}(x_{0})=\psi_{n}=-\frac{2\sum_{s,q=1}^{n-1}w_{s}w_{sq}\beta_{q}}{\beta_{\nu}}-2\sum_{s,q=1}^{n-1}w_{s}w_{q}\left(\frac{\beta_{q}}{\beta_{\nu}}\right)_{s}-2\sum_{s,q=1}^{n-1}\mathcal{K}_{sq}w_{s}w_{q}.$
According to the fact that $x_{0}$ is the maximum point of $\widetilde{G}(x)$ on the boundary, we have
(3.10) $0=\widetilde{G}_{s}(x_{0})=\frac{\psi_{s}}{\psi}+g^{\prime}u_{s}+\alpha_{0}d_{s}=\frac{\psi_{s}}{\psi}+g^{\prime}u_{s},\quad\text{for}\;s=1,2,\cdots,n-1,$
and
(3.11) $0\geq\widetilde{G}_{\nu}(x_{0})=\widetilde{G}_{n}=\frac{\psi_{n}}{\psi}+g^{\prime}u_{n}+\alpha_{0}.$
By (3.9) and (3.10), we obtain
(3.12) $0\geq\widetilde{G}_{n}=$ $\;-\frac{2w_{n}\sum_{s,q=1}^{n-1}w_{s}\beta_{q}d_{sq}}{\beta_{\nu}\psi}-\frac{2\sum_{s,q=1}^{n-1}w_{s}w_{q}\left(\frac{\beta_{q}}{\beta_{\nu}}\right)_{p}}{\psi}$
$\;-\frac{2\sum_{s,q=1}^{n-1}\mathcal{K}_{sq}w_{s}w_{q}}{\psi}+\frac{g^{\prime}\varphi}{\cos\theta}+\alpha_{0}$
$=$ $\;-\frac{2\sum_{i,k=1}^{n-1}\kappa_{i}w_{i}w_{k}\beta_{i}\beta_{k}}{\beta_{\nu}^{2}\sum_{i,k=1}^{n}C^{ik}w_{i}w_{k}}-\frac{2\sum_{i,k=1}^{n-1}w_{i}w_{k}\left(\frac{\beta_{k}}{\beta_{\nu}}\right)_{i}}{\sum_{i,k=1}^{n}C^{ik}w_{i}w_{k}}$
$\;-\frac{2\sum_{s,q=1}^{n-1}\mathcal{K}_{sq}w_{s}w_{q}}{\sum_{i,k=1}^{n}C^{ik}w_{i}w_{k}}+\frac{g^{\prime}\varphi}{\cos\theta}+\alpha_{0}$
$\geq$ $\;-C+\alpha_{0}>0,$
where the last inequality comes from taking $\alpha_{0}$ large enough depending on $|\beta|_{C^{1}}$, $n$, $\beta_{0}$ and the geometry of $\partial\Omega$. Thus this case does not occur at all.
$\mathbf{Case\;\mathrm{\mathbf{II}}}$: $x_{0}\in\Omega_{\mu}$. Assume that $|Du|$ attains its boundary maximum at $y_{0}\in\partial\Omega$ and
(3.13) $|Du|^{2}(y_{0})\geq 4\sup_{\partial\Omega}\left|\frac{\varphi}{\cos\theta}\right|^{2}.$
Since $\widetilde{G}(x_{0})\geq\widetilde{G}(y_{0})$,
(3.14) $\psi(x_{0})\geq\frac{\beta_{0}^{2}C(\mu,\alpha_{0})\text{e}^{-\frac{1}{4}}}{4}|Du|^{2}(y_{0}).$
Then
(3.15) $\psi(x_{0})\geq\frac{\beta_{0}^{2}C(\mu,\alpha_{0})}{8C\text{e}}|Du|^{2}(x_{0})\geq\frac{\beta_{0}^{2}C(\mu,\alpha_{0})}{16C\text{e}}|Dw|^{2}(x_{0}):=C_{0}|Dw|^{2}(x_{0}),$
where denotes $C_{0}\in(0,1)$. Choosing the coordinate such that matrix $D^{2}u(x_{0})$ is diagonal. For $k=1,2,\cdots,n$, denote $T_{k}=\sum_{l=1}^{n}C^{kl}w_{l}$ and $\overrightarrow{T}=(T_{1},T_{2},\cdots,T_{n})$, it is direct to observe that $\left|\overrightarrow{T}\right|\leq|Dw|$ and $\psi\leq\left|\overrightarrow{T}\right||Dw|$. Then
(3.16) $C_{0}|Dw|\leq\left|\overrightarrow{T}\right|\leq|Dw|.$
Without loss of generality, we further assume that $T_{1}w_{1}\geq\frac{C_{0}}{n}|Dw|^{2}>0$ and therefore we have $\frac{w_{1}}{T_{1}}\geq\frac{C_{0}}{n}$. Let $\mu$ is small such that
(3.17) $\frac{u_{1}}{T_{1}}\geq\frac{C_{0}}{3n}.$
At $x_{0}$,
(3.18) $0=\widetilde{G}_{i}=\frac{\psi_{i}}{\psi}+g^{\prime}u_{i}+\alpha_{1}d_{i},$
denote ${C^{kl}}_{,i}=\frac{\partial C^{kl}}{\partial x_{i}}$ and ${C^{kl}}_{,ij}=\frac{\partial^{2}C^{kl}}{\partial x_{i}\partial x_{j}}$, especially for $i=1$ in equation (3.18), we obtain
(3.19) $\sum_{l=1}^{n}T_{l}w_{l1}=-\frac{\psi}{2}(g^{\prime}u_{1}+\alpha_{1}d_{1})-\sum_{k,l=1}^{n}\frac{{C^{kl}}_{,1}}{2}w_{k}w_{l}.$
Take $\mu$ small enough and $|Dw|$ large enough, then
(3.20) $u_{11}\left(1-\varphi_{z}\frac{d}{\cos\theta}\right)<0.$
Now, it is turn to deal with second order derivatives of $\widetilde{G}$. Using the fact that $F$ is homogeneous of degree $(k-l)$, we have
(3.21) $0\geq F^{ij}\widetilde{G}_{ij}(x_{0})=&\;\frac{F^{ij}\left(\sum_{k,l=1}^{n}C^{kl}w_{k}w_{l}\right)_{ij}}{\psi}-2\alpha_{1}g^{\prime}F^{ij}u_{i}d_{j}-\alpha_{1}^{2}F^{ij}d_{i}d_{j}\\ &\;+\left[g^{\prime\prime}-\left(g^{\prime}\right)^{2}\right]F^{ij}u_{i}u_{j}+g^{\prime}F^{ij}u_{ij}+\alpha_{1}F^{ij}d_{ij}\\ \geq&\;\frac{F^{ij}\left(\sum_{k,l=1}^{n}C^{kl}w_{k}w_{l}\right)_{ij}}{\psi}-2\alpha_{1}^{2}F^{ij}d_{i}d_{j}+\alpha_{1}F^{ij}d_{ij}\\ &\;+\left[g^{\prime\prime}-2\left(g^{\prime}\right)^{2}\right]F^{ij}u_{i}u_{j}+g^{\prime}(k-l)f\\ :=&\;\text{I}+\text{II}+\text{III}+\text{IV}+\text{V}.\end{aligned}$
According to (3.16) and (3.17), we obtain
[…]
where $K_{0}$ is a positive constant related to the geometry of $\partial\Omega$.
Next, we just need to deal with the term I
[…]
$=$ $\;\frac{\sum_{i,j,k,l=1}^{n}F^{ij}{C^{kl}}_{,ij}w_{k}w_{l}}{\psi}+\frac{2\sum_{i,j,k,l=1}^{n}F^{ij}C^{kl}w_{ijk}w_{l}}{\psi}$
$\;\frac{4\sum_{i,j,k,l=1}^{n}F^{ij}{C^{kl}}_{,j}w_{ik}w_{l}}{\psi}+\frac{2\sum_{i,j,k,l=1}^{n}F^{ij}C^{kl}w_{ik}w_{jl}}{\psi}$
::=:= $\;I_{1}+I_{2}+I_{3}+I_{4}.$
For the term $I_{1}$, it is easy to know that
(3.24) $\psi I_{1}=\sum_{i,j,k,l=1}^{n}F^{ij}{C^{kl}}_{,ij}w_{k}w_{l}\geq-C|Dw|^{2}\sum_{i=1}^{n}F^{ii}.$
For the term $I_{2}$, by (2.4), (3.16) and Lemma 2.5, we derive
(3.25) $\psi I_{2}=$  $\;2\sum_{k=1}^{n}T_{k}\left[\partial_{k}f-\sum_{i,j=1}^{n}F^{ij}\left(\frac{\varphi d}{\cos\theta}\right)_{ijk}\right]$
$\geq$ $\;-Cd|Dw|^{4}\sum_{i=1}^{n}F^{ii}-C|Dw|^{3}\sum_{i=1}^{n}F^{ii}$
$\;-Cd|Dw|^{2}\sum_{i=1}^{n}\left|F^{ii}u_{ii}\right|-C|Dw|\sum_{i=1}^{n}\left|F^{ii}u_{ii}\right|.$
Similarly,
(3.26) $\psi I_{3}=4\sum_{i,j,k,l=1}^{n}F^{ij}{C^{kl}}_{,j}w_{ik}w_{l}\geq-C|Dw|^{3}\sum_{i=1}^{n}F^{ii}-C|Dw|\sum_{i=1}^{n}\left|F^{ii}u_{ii}\right|,$
(3.27) $\psi I_{4}=2\sum_{i,j,k,l=1}^{n}F^{ij}C^{kl}w_{ik}w_{jl}\geq$  $\;2\sum_{i=1}^{n}F^{ii}C^{ii}u_{ii}^{2}-Cd|Dw|^{2}\sum_{i=1}^{n}\left|F^{ii}u_{ii}\right|$
$\;-Cd|Dw|^{4}\sum_{i=1}^{n}F^{ii}-C|Dw|^{3}\sum_{i=1}^{n}F^{ii}.$
Combining (3.23)-(3.27), we obtain
(3.28) $\psi I\geq$  $\;2\sum_{i=1}^{n}F^{ii}C^{ii}u_{ii}^{2}-Cd|Dw|^{2}\sum_{i=1}^{n}\left|F^{ii}u_{ii}\right|-C|Dw|\sum_{i=1}^{n}\left|F^{ii}u_{ii}\right|$
$\;-Cd|Dw|^{4}\sum_{i=1}^{n}F^{ii}-C|Dw|^{3}\sum_{i=1}^{n}F^{ii}.$
Let
(3.29) $\mathcal{H}=2\sum_{i=1}^{n}F^{ii}C^{ii}u_{ii}^{2}-Cd|Dw|^{2}\sum_{i=1}^{n}\left|F^{ii}u_{ii}\right|-C|Dw|\sum_{i=1}^{n}\left|F^{ii}u_{ii}\right|,$
and we will bound $\mathcal{H}$ from below in the following.
Without loss of generality, we can set $C^{11}$ be the smallest of $\left\{C^{ii}\right\}_{i=1}^{n}$. Then, we have $C^{ii}\geq\frac{1}{2}$ for any $i\geq 2$. otherwise it follows that $\sum_{i=1}^{n}C^{ii}<n-1$, which contradicts with $\sum_{i=1}^{n}C^{ii}=n-1$. A simple calculation yields
(3.30) $\mathcal{H}\geq-C\sum_{i=1}^{n}F^{ii}\left(d|Dw|^{2}+|Dw|\right)^{2}.$
Plugging (3.30) into (3.28), then
(3.31) $I\geq-Cd|Dw|^{2}\sum_{i=1}^{n}F^{ii}-C|Dw|\sum_{i=1}^{n}F^{ii}.$
Substituting (3.22) and (3.31) into (3.21), we arrive at
(3.32) $0\geq\left\{C_{2}\left[g^{\prime\prime}-2\left(g^{\prime}\right)^{2}\right]|Dw|^{2}-C_{3}\mu|Dw|^{2}-C|Dw|\right\}\sum_{i=1}^{n}F^{ii}.$
Let $\mu$ be small enough such that $C_{2}\left(\frac{1}{16M}\right)^{2}|Dw|^{2}-C|Dw|\leq 0$, then the proof of Theorem 3.1 have been finished.∎
4. Second order estimates
4.1. Global second order estimates
In this subsection, we derive global second order estimates for admissible solutions of problem (1.1).
Lemma 4.1.
Let $u\in C^{4}(\Omega)\cap C^{3}(\overline{\Omega})$ be an admissible solution of problem (1.1) in a bounded $C^{3}$ domain $\Omega\subset\mathbb{R}^{n}$. Suppose $f^{\frac{1}{k-l}}(x,z)\in C^{1,1}(\overline{\Omega}\times\mathbb{R})$, then there exists a positive constant $C$ depending on $|f^{\frac{1}{k-l}}|_{C^{1,1}}$, $\overline{\Omega}$, $n$, $k$, $l$, $\gamma$ such that
(4.1) $\sup_{\Omega}|D^{2}u|\leq\sup_{\partial\Omega}|D^{2}u|+C.$
Proof.
Consider an auxiliary function for any $\xi=(\xi^{1},\xi^{2},\cdots,\xi^{n})\in\mathbb{R}^{n}$
$V(x,\xi)=u_{\xi\xi}+\frac{1}{2}u_{\xi}^{2}.$
Assume that $V$ attains its the interior maximum at $\left(x_{0},\xi_{0}\right)$. We may also suppose $\xi_{0}=(1,0,\cdots,0)$ and the matrix $u_{ij}(x_{0})$ is diagonal with $u_{11}(x_{0})\geq u_{22}(x_{0})\geq\cdots\geq u_{nn}(x_{0})$ under an orthogonal argument. Set
(4.2) $\widetilde{V}=\;u_{11}+\frac{1}{2}u_{1}^{2},$
it follows that, at $x_{0}$
(4.3) 00 0 $=\widetilde{V}_{i}=u_{11i}+u_{1}u_{1i},$
(4.4) 00 0 $\geq\widetilde{V}_{ij}=u_{11ij}+u_{1i}u_{1j}+u_{1}u_{1ij}.$
In the rest of the proof, all the calculations are made at $x_{0}$. Since $f^{\frac{1}{k-l}}\in C^{1,1}$ and Lemma 2.5, we obtain
(4.5) $\partial_{11}f\geq-C\sum_{i=1}^{n}F^{ii}+\left(1-\frac{1}{k-l}\right)\frac{(\partial_{1}f)^{2}}{f}.$
By (2.4)-(2.6), (4.4)-(4.5) and Lemma 2.4
(4.6) $\begin{split}0\geq L\widetilde{V}=&\;F^{ij}u_{11ij}+F^{ij}u_{1i}u_{1j}+F^{ij}u_{1}u_{1ij}\\ \geq&\;-C\sum_{i=1}^{n}F^{ii}+u_{11}^{2}\sum_{i=1}^{n}F^{ii},\end{split}$
which means that
(4.7) $\sup_{\Omega}|D^{2}u|\leq\sup_{\partial\Omega}|D^{2}u|+C.$
4.2. Boundary second order estimates
In this part, we establish boundary second order estimates, and divide the proof into three cases: double tangential derivatives, mixed tangential-oblique derivatives and double oblique derivatives. We assume that the function $G(x,z,p)$ and the unit inner normal vector field $\nu$ have been extended to $\overline{\Omega}$, to be constant along normals to $\partial\Omega$ in some neighbourhood $\Omega_{0}$ of $\partial\Omega$.
Lemma 4.2.
Let $u\in C^{4}(\Omega)\cap C^{3}(\overline{\Omega})$ be an admissible solution of problem (1.1) in a bounded $C^{3}$ domain $\Omega\subset\mathbb{R}^{n}$. Then for any tangential vector field $\tau$, we have
(4.8) $\sup_{\partial\Omega}|u_{\tau\beta}|\leq C,$
where $\beta=G_{p}(x,z,p)$ and the positive constant $C$ depends on $G,\overline{\Omega},|u|_{C^{1}}$.
Proof.
Differentiating $G(x,z,p)=0$ in problem (1.1) with respect to $\tau$, we have
(4.9) $D_{\tau}G=\left\langle DG,\tau\right\rangle=G_{x_{\tau}}+G_{z}u_{\tau}+G_{p_{k}}u_{k\tau}=0.$
Hence, we get $|u_{\tau\beta}|\leq C$ on $\partial\Omega$. ∎
For convenience, we set
$M_{2}=\sup_{\Omega}\left|D^{2}u\right|,\quad M_{2}^{\prime}=\sup_{\partial\Omega}\sup_{|\tau|=1,\left\langle\tau,\nu\right\rangle=0}\left|u_{\tau\tau}\right|.$
Next, we estimate double tangential derivatives on the boundary.
Lemma 4.3.
Let $u\in C^{4}(\Omega)\cap C^{3}(\overline{\Omega})$ be an admissible solution of problem (1.1) in a bounded $C^{3}$ domain $\Omega\subset\mathbb{R}^{n}$. Suppose that $G\in C^{2}(\partial\Omega\times\mathbb{R}\times\mathbb{R}^{n})$ satisfy the boundary condition (1.2) with $G_{pp}\leq 0$ and $f^{\frac{1}{k-l}}(x,z)\in C^{1,1}(\overline{\Omega}\times\mathbb{R})$. Then for any tangential vector field $\tau$ with $|\tau|\leq 1$ and constant $\varepsilon>0$, we have
(4.10) $\sup_{\partial\Omega}u_{\tau\tau}\leq\varepsilon M_{2}+C_{\varepsilon},$
where $C_{\varepsilon}$ is a positive constant depending on $\varepsilon$, $\overline{\Omega}$, $G$, $n$, $k$, $l$, $\gamma$, $|f^{\frac{1}{k-l}}|_{C^{1,1}}$, $\beta_{0}$, $|u|_{C^{1}}$.
Proof.
Suppose a function
(4.11) $v_{\tau}=u_{\tau\tau}+\frac{1}{2}|u_{\tau}|^{2}$
attains a maximum over $\partial\Omega$ and tangential vector $\tau$ satisfied $|\tau|\leq 1$, at $y_{0}\in\partial\Omega$ and vector $\tau=\tau_{0}$. Assume $y_{0}=0$ and $\tau_{0}=e_{1}=(1,0,\cdots,0)$. Set $b=\frac{\nu_{1}}{\left\langle\beta,\nu\right\rangle}$ and $\tau=e_{1}-b\beta$, then, at any point on $\partial\Omega$
(4.12) $v_{1}=v_{\tau}+b\left(2u_{\tau\beta}+u_{\beta}u_{\tau}\right)+b^{2}\left(u_{\beta\beta}+\frac{1}{2}u_{\beta}^{2}\right)$
satisfy $v_{1}(0)=v_{\tau}(0)$, $b(0)=0$ and $\tau(0)=e_{1}$. Let $h=\frac{1}{\left\langle\beta,\nu\right\rangle}(2u_{\tau\beta}+u_{\beta}u_{\tau})$, then
(4.13) $\left|h-h(0)\right|\leq C(1+M_{2})|x|,\quad\text{on}\;\partial\Omega,$
and
(4.14) $v_{1}-h(0)\nu_{1}\leq v_{\tau}+C_{1}(1+M_{2})|x|^{2},$
where $C_{1}$ is a positive constant depending on $G$, $\beta_{0}$, $\overline{\Omega}$, $|u|_{C^{1}}$. Thus, we set
(4.15) $\widetilde{v}_{1}:=v_{1}-h(0)\nu_{1}-C_{1}(1+M_{2})|x|^{2},$
where the constant $C_{1}$ is the same quantities in (4.14) satisfying
(4.16) $\widetilde{v}_{1}\leq\;v_{\tau}=u_{\tau\tau}+\frac{1}{2}|u_{\tau}|^{2}\leq|\tau|^{2}v_{1}(0)\leq Qv_{1}(0),$
here $Q$ is a non-negative function in $C^{2}(\overline{\Omega})$ satisfying $Q\geq|\tau|^{2}$ on $\partial\Omega$ and $Q(0)=1$. Remark that
[…]
Assume $v_{1}(0)>0$ and $C_{1}$ large enough such that $Q>0$ in $\overline{\Omega}$. Now differentiating $G(x,z,p)=0$ twice in direction $\tau$ with $\tau(0)=e_{1}$ and using $G_{pp}\leq 0$, we obtain
(4.18) $D_{\beta}u_{11}(0)\geq\;-C\left(1+M_{2}\right).$
Next, we consider a new function
(4.19) $\overline{v}=\widetilde{v}_{1}-v_{1}(0)Q-K(1+M_{2})\phi,$
where $\phi\in C^{2}(\overline{\Omega})$ is a negative define function in $\Omega$ satisfying $\phi=0$ on $\partial\Omega$, $\phi_{\nu}=-1$ and $\phi_{\tau}=0$ on $\partial\Omega$, $K$ is a sufficiently large constant to be determined later. At $y_{0}=0$,
(4.20) $\begin{split}D_{\beta}\overline{v}=&\;D_{\beta}\left[\widetilde{v}_{1}-v_{1}(0)Q\right]-K(1+M_{2})\phi_{\beta}>0\end{split}$
with $K>\frac{C_{1}}{\beta_{0}}$. The specific details can be referred to Jiang-Trudinger [21, 23].
Since $\phi=0$ on $\partial\Omega$ and $\widetilde{v}_{1}\leq v_{1}(0)Q$, then $\overline{v}\leq 0\;\text{on}\;\partial\Omega$ and $\overline{v}(0)=0$. Hence $\overline{v}$ takes its maximum at $\overline{y}_{0}\in\Omega$. We suppose that the matrix $u_{ij}(\overline{y}_{0})$ is diagonal. According to (4.5) and Lemma 2.5, we have
(4.21) $0\geq L\overline{v}(\overline{y}_{0})=$ $\;F^{ij}u_{11ij}+F^{ij}u_{1i}u_{1j}+u_{1}F^{ij}u_{1ij}$
$\;-h(0)F^{ij}\nu_{1ij}-2C_{1}(1+M_{2})\sum_{i=1}^{n}F^{ii}$
$\geq$ $\;-C(1+M_{2})\sum_{i=1}^{n}F^{ii}+Cu_{11}^{2}\sum_{i=1}^{n}F^{ii},$
then $u_{11}(\overline{y}_{0})\leq\sqrt{C(1+M_{2})}\leq\varepsilon M_{2}+C_{\varepsilon}$, where $\varepsilon$ is any positive constant and $C_{\varepsilon}$ depends on $\varepsilon$, $n$, $k$, $l$, $\gamma$, $\overline{\Omega}$, $G$, $|f^{\frac{1}{k-l}}|_{C^{1,1}}$, $\beta_{0}$, $|u|_{C^{1}}$. Set $\phi\geq-\frac{\varepsilon}{4K}$ and $Q>\frac{1}{2}$, we get
(4.22) $v_{1}(0)\leq\varepsilon M_{2}+C_{\varepsilon}.$
Since $v_{\tau}$ attains the maximum at 00 and $\tau=e_{1}$ on $\partial\Omega$, then $v_{\tau}\leq v_{1}(0)\leq\varepsilon M_{2}+C_{\varepsilon}$, which implies that
(4.23) $\sup_{\partial\Omega}u_{\tau\tau}\leq\varepsilon M_{2}+C_{\varepsilon}.$
Lemma 4.4.
Let $u\in C^{4}(\Omega)\cap C^{3}(\overline{\Omega})$ be an admissible solution of problem (1.1) in a bounded $C^{3}$ domain $\Omega\subset\mathbb{R}^{n}$. Suppose that $G\in C^{2}(\partial\Omega\times\mathbb{R}\times\mathbb{R}^{n})$ satisfy the boundary condition (1.2) with $G_{pp}=0$ and $f^{\frac{1}{k-l}}(x,z)\in C^{1,1}(\overline{\Omega}\times\mathbb{R})$. Then for any $\varepsilon>0$,
(4.24) $\sup_{\partial\Omega}u_{\beta\beta}\leq\varepsilon M_{2}+C_{\varepsilon}\left(1+M_{2}^{\prime}\right),$
where $C_{\varepsilon}$ is a positive constant depending on $\varepsilon$, $\overline{\Omega}$, $G$, $n$, $k$, $l$, $\gamma$, $|f^{\frac{1}{k-l}}|_{C^{1,1}}$, $\beta_{0}$, $|u|_{C^{1}}$.
Proof.
For any fixed boundary point $x_{0}\in\partial\Omega$, we consider the function
(4.25) $\overline{V}=G(x,z,p)+\frac{a_{1}}{2}|Du-Du(x_{0})|^{2},$
here $a_{1}\leq 1$ is a positive constant. Since $\widetilde{F}$ is homogeneous of degree $1$ and $G_{pp}=0$, we have
(4.26) $\widetilde{L}\overline{V}=$ $\;\widetilde{F}^{ij}G_{x_{i}x_{j}}+2\widetilde{F}^{ij}G_{x_{i}z}u_{j}+2\widetilde{F}^{ij}G_{x_{i}p_{k}}u_{jk}+G_{zz}\widetilde{F}^{ij}u_{i}u_{j}$
$\;+2\widetilde{F}^{ij}G_{zp_{k}}u_{i}u_{jk}+G_{z}\widetilde{F}^{ij}u_{ij}+G_{p_{k}p_{l}}\widetilde{F}^{ij}u_{ik}u_{lk}+G_{p_{k}}\widetilde{F}^{ij}u_{ijk}$
$\;+a_{1}\sum_{k=1}^{n}\widetilde{F}^{ij}u_{ik}u_{jk}+a_{1}\sum_{k=1}^{n}(u_{k}-u_{k}(x_{0}))\widetilde{F}^{ij}u_{ijk}$
$\geq$ $\;-C\sum_{i=1}^{n}\widetilde{F}^{ii}+a_{1}\sum_{k=1}^{n}\widetilde{F}^{ij}u_{ik}u_{jk}+2\widetilde{F}^{ij}G_{x_{i}p_{k}}u_{jk}+2\widetilde{F}^{ij}G_{zp_{k}}u_{i}u_{jk}$
$\geq$ $\;-C\left(1+\frac{1}{a_{1}}\right)\sum_{i=1}^{n}\widetilde{F}^{ii},$
where $\widetilde{L}=\widetilde{F}^{ij}\partial_{ij}$ and the last inequality used $\sum_{i=1}^{n}\widetilde{F}^{ii}\geq C(n,k,l,\gamma)>0$.
Now, a suitable upper barrier for $\overline{V}$ at the point $x_{0}\in\partial\Omega$ can be constructed as
(4.27) $\overline{\phi}=-c\phi+\frac{a_{2}}{2}|x-x_{0}|^{2},\quad\quad\text{in}\;\Omega_{\mu},$
with
(4.28) $\phi=-d+K_{3}d^{2},$
[…]
(4.29) $\widetilde{L}\overline{\phi}=-c\widetilde{F}^{ij}\phi_{ij}+a_{2}\widetilde{F}^{ij}\delta_{ij}\leq(a_{2}-c\delta)\sum_{i=1}^{n}\widetilde{F}^{ii},\quad\text{in}\;\Omega_{\mu},$
provided $\mu\leq\frac{1}{4K_{3}}$, where $\delta>0$ small enough and $K_{3}$ large enough depending on the curvature of $\partial\Omega$, $n$, $k$, $l$. Let $c\geq\frac{a_{2}}{\delta}+\frac{C}{\delta}\left(1+\frac{1}{a_{1}}\right)$, we obtain
(4.30) $\widetilde{L}\left(\overline{V}-\overline{\phi}\right)\geq 0,\quad\text{in}\;\Omega_{\mu}.$
Next, we deal with $\overline{V}$ and $\overline{\phi}$ on the boundary of $\Omega_{\mu}$.
$\mathbf{Case\;\mathrm{\mathbf{I}}}$: For any $x\in\partial\Omega$, $\overline{V}=\frac{a_{1}}{2}|Du-Du(x_{0})|^{2}$ and $\overline{\phi}=\frac{a_{2}}{2}|x-x_{0}|^{2}$, then
(4.31) $\begin{split}|Du-Du(x_{0})|\leq C(1+M_{2}^{\prime})|x-x_{0}|.\end{split}$
For any $y\in\Omega_{\mu}$ and $y^{\prime}$ the closest point on $\partial\Omega$, we obtain
(4.32) $\begin{split}|Du(y)-Du(x_{0})|^{2}\leq C(1+(M_{2}^{\prime})^{2}+M_{2})\left(|y^{\prime}-x_{0}|^{2}+d\right).\end{split}$
Choose $a_{2}=a_{1}(1+(M_{2}^{\prime})^{2})+M_{2})$ and $c\geq a_{2}$ such that
(4.33) $\overline{V}-\overline{\phi}\leq 0,\quad\text{on}\;\overline{\Omega}_{\mu}\cap\left\{G(x,z,p)=0\right\}\supset\partial\Omega.$
$\mathbf{Case\;\mathrm{\mathbf{II}}}$: On $\partial\Omega_{\mu}\cap\Omega$, by choosing $c\geq\frac{C}{\mu}$, we have
(4.34) $\overline{V}-\overline{\phi}\leq 0,\quad\quad\quad\text{on}\;\partial\Omega_{\mu}\cap\Omega.$
Therefore,
(4.35) $\overline{V}-\overline{\phi}\leq 0,\quad\quad\quad\text{in}\;\Omega_{\mu}.$
Since $\overline{V}(x_{0})=\overline{\phi}(x_{0})=0$,
(4.36) $D_{\beta}\overline{V}(x_{0})\leq D_{\beta}\overline{\phi}(x_{0}),$
which implies
(4.37) $u_{\beta\beta}(x_{0})\leq c\sup_{\partial\Omega}\left\langle\nu,\beta\right\rangle+C.$
Fix the constant $c$ such that
(4.38) $c\leq\frac{a_{2}+C(1+\frac{1}{a_{2}})}{\delta}+a_{2}+\frac{C}{\mu}\leq\frac{C}{\delta}\varepsilon_{1}M_{2}+C_{\varepsilon_{1}}(1+M_{2}^{\prime}),$
where $C$ depends on $G$, $n$, $k$, $l$, $\gamma$, $\overline{\Omega}$, $|f^{\frac{1}{k-l}}|_{C^{1,1}}$, $|u|_{C^{1}}$ and $\beta_{0}$. For any $\varepsilon>0$, taking $a_{1}=\frac{1}{1+\varepsilon_{1}M_{2}}$ and $\varepsilon_{1}\leq\frac{\varepsilon\delta}{C\sup_{\partial\Omega}\left\langle\nu,\beta\right\rangle}$ for a further constant $C$ in (4.38), we get
(4.39) $u_{\beta\beta}(x_{0})\leq\varepsilon M_{2}+C_{\varepsilon}(1+M_{2}^{\prime}).$
Thus, Lemma 4.4 can be obtained by the arbitrariness of $x_{0}$ and Lemma 4.3. ∎
We then directly arrive at the following $C^{2}$ estimates.
Theorem 4.5.
Let $u\in C^{4}(\Omega)\cap C^{3}(\overline{\Omega})$ be an admissible solution of the boundary value problem (1.1) in a bounded $C^{3}$ domain $\Omega\subset\mathbb{R}^{n}$. Suppose that $G\in C^{2}(\partial\Omega\times\mathbb{R}\times\mathbb{R}^{n})$ satisfy the boundary condition (1.2) with $G_{pp}=0$ and $f^{\frac{1}{k-l}}(x,z)\in C^{1,1}(\overline{\Omega}\times\mathbb{R})$. Then
(4.40) $\sup_{\overline{\Omega}}|D^{2}u|\leq C,$
where $C$ is a positive constant depending on $\overline{\Omega}$, $G$, $n$, $k$, $l$, $\gamma$, $|f^{\frac{1}{k-l}}|_{C^{1,1}}$, $\beta_{0}$, $|u|_{C^{1}}$.
Proof.
Combining Lemma 4.1-Lemma 4.4, we obtain
(4.41) $\sup_{\partial\Omega}|u_{\xi\xi}|\leq\varepsilon M_{2}+C_{\varepsilon},$
for any constant unit vector $\xi$ and constant $\varepsilon>0$. Applying global second order derivative estimates, we get
(4.42) $\sup_{\overline{\Omega}}|D^{2}u|\leq C.$
Finally, we complete the proof of Theorem 1.3. From the maximal principle, $C^{0}$ estimates can be directly obtained by the existence of the admissible subsolution and $f_{z}\geq 0,\ \varphi_{z}>0$. Then based on the a priori estimates for non-degenerate case, $C^{2,\alpha}\;(0<\alpha<1)$ estimates can be established by Evans-Krylov theory and higher order estimates followed by Schauder theory. Then the existence result can be derived by the continuity method and the uniqueness assertion is immediate from the maximum principle, more details see [15]. Theorem 1.3 can be proved by approximation as in [25].